A charge of 2 µC is 0.5m from a charge of 3µC. Find the electric force.


1
Expert's answer
2021-06-24T16:14:15-0400

"F=k\frac{q_1\cdot q_2}{x^2}=9\cdot10^9\cdot\frac{2\cdot10^{-6}\cdot3\cdot10^{-6}}{0.5^2}=0.216\ (N)" . Answer